# TPP-LLM: Modeling Temporal Point Processes by Efficiently Fine-Tuning Large Language Models

### v1

@@ 6 Conclusion
We have made several efforts to ensure the reproducibility of our findings by providing detailed documentation of the experimental setup and methodologies. Detailed descriptions of the datasets used in our experiments, including the preprocessing steps and sequence structure, can be found in Appendix A. The code used for implementing our TPP-LLM model, including the fine-tuning mechanisms, is made available on the GitHub repository [^2] https://github.com/zefang-liu/TPP-LLM. Model architecture details, training procedures, and hyperparameter settings, are provided in Section 5 and Appendix D, enabling the replication of experiments. Additionally, the theoretical foundations of our model are fully explained in Section 3 and 4. This comprehensive approach is intended to facilitate the reproduction of our results and to support further research building on our work.
@@ A.3 Event Types
In this subsection, we provide a detailed mapping of event types to their corresponding textual descriptions for each dataset in Table 11-15. These event types represent the various categories of events modeled in our experiments, allowing the model to capture diverse patterns across different domains.
@@ 3 Queens Pickup 7 Bronx Pickup
ID Event Type
0 Large
1 Medium
2 Small
@@ 8 Toys and Games 17 Cell Phones and Accessories
Appendix B Baseline Details
Neural Hawkes Process (NHP) (Mei & Eisner, 2017) is a generative model that uses a continuous-time LSTM to dynamically adjust the intensity of multiple event types based on the sequence of past events, enabling accurate predictions of future event types and timings. Self-Attentive Hawkes Process (SAHP) (Zhang et al., 2020) enhances Hawkes processes by using self-attention to model event dynamics, incorporating time intervals into positional encoding, and improving predictive accuracy and interpretability compared to RNN-based models. Transformer Hawkes Process (THP) (Zuo et al., 2020) leverages the self-attention mechanism to efficiently capture long-term dependencies in event sequence data, improving prediction accuracy and likelihood over recurrent neural network-based models. Attentive Hawkes Process (AttNHP) (Yang et al., 2022) replaces LSTM-based architectures with attention-based models to more efficiently capture event sequences and participant embeddings, maintaining or improving prediction accuracy compared to previous neuro-symbolic and attention-based approaches.
Appendix C Prompt Details
@@ Appendix D Model Hyperparameters
Table 17: Hyperparameter configurations used for various models in the experiments. The hidden size, embedding dimension, and other model structure parameters of TPP-LLM depend on the foundation model.
Hyperparameter NHP SAHP THP AttNHP TPP-LLM
hidden_size 64 64 64 64 -
time_emb_size 16 16 16 16 -
num_layers 2 2 2 2 -
num_heads - 2 2 2 -
batch_size 128 128 128 8 8
max_epoch 10 20 20 8 20
learning_rate 1e-3 1e-3 1e-3 1e-3 5e-4
num_integral_samples 20 20 20 20 20
lora_rank - - - - 16
lora_alpha - - - - 16
lora_dropout - - - - 0.05
target_modules - - - - Q, K, V, O
beta_type - - - - 1
beta_time - - - - 1
Appendix E Experiment Result Details
E.1 Standard Deviations for Full-Training Results
This subsection reports the standard deviations for log-likelihood (Table 18), next event type prediction accuracy (Table 19), and next event time prediction RMSE (Table 20) from the full-training experiments. These results help assess the consistency of model performance across different datasets.
Table 18: Performance comparison of standard deviations in log-likelihood across different datasets.
@@ E.1 Standard Deviations for Full-Training Results
NHP 0.0094 0.0134 0.0074 0.0019 0.0236
SAHP 0.0781 0.0140 0.0765 0.0140 0.1167
THP 0.0100 0.0046 0.0794 0.0084 0.0192
AttNHP 0.0095 0.0025 0.0353 0.0070 0.0205
TPP-Llama 0.0137 0.0160 0.0497 0.0075 0.0392
TPP-Gemma 0.0118 0.0098 0.0086 0.0024 0.0380
Table 19: Performance comparison of standard deviations in next event type prediction accuracy across different datasets.
@@
NHP 0.0068 0.0066 0.0015 0.0019 0.0056
SAHP 0.0249 0.0075 0.0052 0.0097 0.0066
THP 0.0020 0.0014 0.0034 0.0035 0.0037
AttNHP 0.0129 0.0017 0.0731 0.0039 0.0009
TPP-Llama 0.0014 0.0033 0.0039 0.0028 0.0021
TPP-Gemma 0.0013 0.0058 0.0021 0.0030 0.0050
Table 20: Performance comparison of standard deviations in next event time prediction RMSE across different datasets.
Model StackOverflow Crime Taxi Earthquake Amazon
NHP 0.0048 0.0052 0.0075 0.0105 0.0059
SAHP 0.0003 0.0005 0.0021 0.0053 0.0004
THP 0.0056 0.0065 0.0179 0.0024 0.0153
AttNHP 0.0079 0.0021 0.0054 0.0089 0.0024
TPP-Llama 0.0059 0.0053 0.0490 0.0061 0.0018
TPP-Gemma 0.0047 0.0116 0.0065 0.0063 0.0065
E.2 Standard Deviations for Few-Shot Results
This subsection presents the standard deviations for log-likelihood (Table 21), accuracy (Table 22), and RMSE (Table 23) from the few-shot learning experiments. These values provide insight into how stable the models are when trained on limited data.
@@ E.2 Standard Deviations for Few-Shot Results
Table 21: Performance comparison of standard deviations in log-likelihood performance across different datasets on the 2% training set.
Model StackOverflow Crime Taxi Earthquake Amazon
NHP 0.1837 0.0899 0.0299 0.0247 0.2063
SAHP 0.0083 0.0061 0.0098 0.0252 0.0042
THP 0.1024 0.0935 0.3683 0.0288 0.0581
AttNHP 0.0254 0.0056 0.0050 0.0173 0.0231
TPP-Llama 0.0107 0.0367 0.2542 0.0173 0.0889
TPP-Gemma 0.0305 0.0283 0.1136 0.0624 0.1998
Table 22: Performance comparison of standard deviations in next event type prediction accuracy across different datasets on the 2% training set.
Model StackOverflow Crime Taxi Earthquake Amazon
NHP 0.0181 0.0636 0.0169 0.0061 0.0069
SAHP 0.0189 0.0728 0.0039 0.0277 0.0003
THP 0.0117 0.0048 0.0002 0.0716 0.0000
AttNHP 0.0014 0.0020 0.0136 0.0077 0.0001
TPP-Llama 0.0022 0.0156 0.0069 0.0036 0.0055
TPP-Gemma 0.0046 0.0101 0.0033 0.0084 0.0055
Table 23: Performance comparison of standard deviations in next event time prediction RMSE across different datasets on the 2% training set.
Model StackOverflow Crime Taxi Earthquake Amazon
NHP 0.0012 0.0007 0.0012 0.0265 0.0017
SAHP 0.0005 0.0005 0.0007 0.0196 0.0009
THP 0.1179 0.0806 0.0032 0.0215 0.0602
AttNHP 0.0121 0.0140 0.0042 0.0221 0.0400
TPP-Llama 0.0268 0.0497 0.0942 0.1706 0.1026
TPP-Gemma 0.0211 0.0149 0.1693 0.1511 0.0346
E.3 Standard Deviations for Fine-Tuning Results
This section provides a detailed comparison of standard deviations for log-likelihood (Table 24), event type prediction accuracy (Table 25), and event time RMSE (Table 26) across different fine-tuning settings. Tables below highlight the performance across datasets when using various LoRA ranks and no fine-tuning.
Table 24: Performance comparison of standard deviations in log-likelihood performance across different datasets with various fine-tuning settings.
Fine-Tuning StackOverflow Earthquake
None 0.0292 0.0138
LoRA (rank 4) 0.0131 0.0049
LoRA (rank 8) 0.0059 0.0087
LoRA (rank 16) 0.0137 0.0075
LoRA (rank 32) 0.0079 0.0055
@@ E.3 Standard Deviations for Fine-Tuning Results
Table 25: Performance comparison of standard deviations in next event type prediction accuracy across different datasets with various fine-tuning settings.
Fine-Tuning StackOverflow Earthquake
None 0.0019 0.0058
LoRA (rank 4) 0.0018 0.0020
LoRA (rank 8) 0.0024 0.0029
LoRA (rank 16) 0.0014 0.0028
LoRA (rank 32) 0.0012 0.0028
Table 26: Performance comparison of standard deviations in next event time prediction RMSE across different datasets with various fine-tuning settings.
Fine-Tuning StackOverflow Earthquake
None 0.0270 0.0185
LoRA (rank 4) 0.0091 0.0082
LoRA (rank 8) 0.0122 0.0220
LoRA (rank 16) 0.0059 0.0061
LoRA (rank 32) 0.0102 0.0076

### v2

@@ 6 Conclusion
We have made several efforts to ensure the reproducibility of our findings by providing detailed documentation of the experimental setup and methodologies. Detailed descriptions of the datasets used in our experiments, including the preprocessing steps and sequence structure, can be found in Section 5.1 and Appendix A. The code used for implementing our TPP-LLM model, including the fine-tuning mechanisms, is made available on the GitHub repository https://github.com/zefang-liu/TPP-LLM. Model architecture details, training procedures, and hyperparameter settings, are provided in Section 5.4, Appendix D, and Appendix E, enabling the replication of experiments. Additionally, the theoretical foundations of our model are fully explained in Section 3 and 4. This comprehensive approach is intended to facilitate the reproduction of our results and to support further research building on our work.
@@ A.3 Event Types
In this subsection, we provide a detailed mapping of event types to their corresponding textual descriptions for each dataset in Table 8-12. These event types represent the various categories of events modeled in our experiments, allowing the model to capture diverse patterns across different domains.
@@ 3 Queens Pickup 7 Bronx Pickup
ID Event Type ID Event Type ID Event Type
0 Large 1 Medium 2 Small
@@ 8 Toys and Games 17 Cell Phones and Accessories
Appendix B Prompt Details
@@
Appendix C Baseline Details
Neural Hawkes Process (NHP) (Mei & Eisner, 2017) is a generative model that uses a continuous-time LSTM to dynamically adjust the intensity of multiple event types based on the sequence of past events, enabling accurate predictions of future event types and timings. Self-Attentive Hawkes Process (SAHP) (Zhang et al., 2020) enhances Hawkes processes by using self-attention to model event dynamics, incorporating time intervals into positional encoding, and improving predictive accuracy and interpretability compared to RNN-based models. Transformer Hawkes Process (THP) (Zuo et al., 2020) leverages the self-attention mechanism to efficiently capture long-term dependencies in event sequence data, improving prediction accuracy and likelihood over recurrent neural network-based models. Attentive Hawkes Process (AttNHP) (Yang et al., 2022) replaces LSTM-based architectures with attention-based models to more efficiently capture event sequences and participant embeddings, maintaining or improving prediction accuracy compared to previous neuro-symbolic and attention-based approaches. ODE-based Temporal Point Process (ODETPP) (Chen et al., 2020) leverages Neural ODEs to model temporal point processes, enabling flexible and high-fidelity representations of event sequences in continuous time by using continuous-time neural networks to condition on event history.
@@ Appendix D Model Hyperparameters
Table 14: Hyperparameter configurations used for various models in the experiments. The model structure parameters of TPP-LLM depend on the foundation model (TinyLlama-1.1B in this table).
Hyperparameter NHP SAHP THP AttNHP ODETPP TPP-LLM
hidden_size 64 64 64 64 64 2048
time_emb_size 16 16 16 16 16 2048
num_layers 2 2 2 2 2 22
num_heads - 2 2 2 - 32
batch_size 128 128 128 8 128 8
max_epoch 10 20 20 8 100 20
learning_rate 1e-3 1e-3 1e-3 1e-3 2e-4 5e-4
num_integrals 20 20 20 20 20 20
lora_rank - - - - - 16
lora_alpha - - - - - 16
lora_dropout - - - - - 0.05
target_modules - - - - - QKVO
beta_type - - - - - 1
beta_time - - - - - 1
Appendix E Experimental Setup Details
For the implementation of temporal point processes (TPPs), we used the EasyTPP framework (Xue et al., 2024) with the PyTorch back-end (Paszke et al., 2019). The large language models (LLMs) were implemented using Hugging Face’s Transformers library (Wolf et al., 2020), and we applied parameter-efficient fine-tuning through the PEFT library (Mangrulkar et al., 2022). To enhance computational efficiency, we employed 4-bit quantization from the bitsandbytes library (Dettmers & Zettlemoyer, 2023).
Appendix F Few-Shot Learning
@@
NHP -6.515 -6.202 -1.236 -0.697 -4.106
SAHP -6.839 -6.301 -1.173 -0.615 -4.567
THP -2.683 -2.926 -0.878 -0.625 -1.768
AttNHP -2.070 -2.594 -0.670 -0.521 -1.474
ODETPP -3.517 -4.370 -1.702 -0.677 -4.539
TPP-Llama -2.026 -2.668 -0.098 -0.539 -1.469
TPP-Gemma -2.068 -2.716 0.051 -0.600 -1.630
@@
ODETPP 32.37%/0.863 18.82%/0.793 48.25%/1.426 60.70%/0.828 58.80%/1.650
Appendix G More Ablation Studies
This section presents additional ablation studies, covering data variations, perturbations, event type formats, prompt configurations, and fine-tuning methods.
G.1 Data Variations
We construct two additional variants of the Stack Overflow dataset to evaluate the model’s performance under different data configurations, as shown in Table 17. The longer variant includes 4 years of data (2020-2023) and focuses on users with higher activity levels, specifically those earning 100–200 badges. This results in significantly longer average sequence lengths, increasing from 56 to 132. The larger variant, in contrast, spans 3 years (2021–2023) and selects users with 40–100 badges, increasing the number of sequences from 3,336 to 8,065 and introducing two additional badge types.
Table 17: Statistics overview for different variants of the Stack Overflow dataset.
Dataset # of Types # of Events # of Seq. Avg. Seq. Length Time Unit
Stack Overflow (Original) 25 187,836 3,336 56.31 Month
Stack Overflow (Longer) 25 282,008 2,131 132.34 Month
Stack Overflow (Larger) 27 458,917 8,065 56.90 Month
As shown in Table 18, we evaluate the performance of various models, including two bassline (THP and AttNHP) and our model TPP-Llama, across different variants of the Stack Overflow dataset, which vary in size and sequence length. Despite longer average sequence lengths and larger numbers of sequences in the longer and larger variants, TPP-Llama consistently outperforms the other baseline models in terms of log-likelihood, accuracy, and RMSE. These results demonstrate that TPP-Llama maintains superior performance even as the dataset becomes larger and the average sequence length increases, highlighting the model’s robustness and its ability to effectively handle both larger volumes of data and longer event sequences.
Table 18: Performance comparison of log-likelihood, accuracy, and RMSE across different variants of the Stack Overflow dataset.
Model StackOverflow-Original StackOverflow-Longer StackOverflow-Larger
THP -1.877/43.81%/0.629 -1.861/41.18%/0.561 -2.354/40.74%/0.926
AttNHP -1.798/39.12%/0.581 -1.820/38.45%/0.524 -2.386/38.00%/0.803
TPP-Llama -1.$7$1.699/42.23%/0.432 -2.223/42.00%/0.694
G.2 Data Perturbations
We compare the performance of the model across different perturbation ratios for the StackOverflow and Earthquake datasets. The purpose of this comparison is to evaluate the robustness of the model under specific levels of dataset perturbation (1%, 5%, and 10%) and to simulate real-world scenarios where data may be noisy or imperfect. The perturbation dataset is generated by applying a random perturbation to the original event times, where a perturbation value is drawn uniformly from the range $[-1,1]$ and scaled by the specified perturbation ratio and each time interval. The perturbed time is then calculated by adding the perturbation to the original event time, ensuring that the perturbed time is not earlier than the previous perturbed time.
As shown in Table 19, some metrics of TPP-Llama improve with small levels of perturbation, likely due to data augmentation, which helps reduce overfitting and improves generalization. As the perturbation ratio increases further, however, performance begins to degrade, with slight increases in RMSE and minor changes in log-likelihood and accuracy, reflecting the model capability of handling noise in the data. Despite these fluctuations, TPP-Llama consistently demonstrates stable performance and overforms other baselines. These results highlight the model’s resilience and ability to maintain high performance even in the presence of data perturbations.
@@ G.2 Data Perturbations
Table 19: Performance comparison of log-likelihood, accuracy, and RMSE across different dataset perturbations.
Model Perturbation Ratio StackOverflow Earthquake
TPP-Llama None -1.$7$0.475/62.70%/0.288
TPP-Llama 1% -1.775/44.21%/0.498 -0.471/63.04%/0.289
TPP-Llama 5% -1.776/44.17%/0.494 -0.473/63.12%/0.294
TPP-Llama 10% -1.776/44.18%/0.495 -0.470/63.09%/0.293
THP None -1.877/43.81%/0.629 -0.513/62.39%/0.377
AttNHP None -1.798/39.12%/0.581 -0.481/61.87%/0.386
G.3 Event Type Formats
In this ablation study, we investigate the impact of using textual descriptions versus ordinal numbers for event types. The textual input uses the event type itself as the event text, while the ordinal input replaces these descriptions with numerical identifiers. As shown in Table 20, the event time prediction remains relatively consistent between the two settings. However, the event type prediction accuracy significantly drops when ordinal numbers are used, particularly for the Stack Overflow dataset, which features more complex and diverse event types compared to the Earthquake dataset. These results underscore the importance of semantic information in textual descriptions for improving event type prediction accuracy, especially in datasets with high variability in event types.
Table 20: Performance comparison of log-likelihood, accuracy, and RMSE with different event type formats.
Event Type Format StackOverflow Earthquake
Textual -1.$7$0.475/62.70%/0.288
Ordinal -1.772/43.42%/0.480 -0.473/62.54%/0.284
G.4 Prompt Settings
Foundation Model Prompt StackOverflow Earthquake
TinyLlama-1.1B-Intermediate Y -1.$7$0.476/62.65%/0.299
TinyLlama-1.1B-Chat-v1.0 Y -1.$7$0.475/62.70%/0.288
G.5 Fine-Tuning Methods
Table 22 and Figure 4 illustrate the impact of different fine-tuning methods on performance. Without fine-tuning (only training the head layers), the model suffers significant drops in log-likelihood and accuracy, highlighting the need for adapting the pretrained LLM. Fine-tuning with LoRA consistently enhances performance, with higher ranks benefiting more complex tasks and lower ranks offering competitive results with less computational cost. Additionally, alternative methods like LoHa (Hyeon-Woo et al., 2022), LoKr (Yeh et al., 2023), and IA3 (Liu et al., 2022) demonstrate unique strengths: LoKr achieves excellent efficiency with the lowest trainable parameters, LoHa shows strong log-likelihood on Stack Overflow, and IA3 performs well on RMSE for event type prediction. These results highlight trade-offs between computational efficiency and predictive performance among fine-tuning methods.
@@ G.5 Fine-Tuning Methods
Table 22: Performance comparison of log-likelihood, next event type prediction accuracy, and next event time prediction RMSE across different datasets with various fine-tuning settings. (Trainable %: Percentages of trainable parameters in the foundation model, excluding prediction head layers.)
Fine-Tuning Quantization Trainable StackOverflow Earthquake
None 4-bit - -1.891/42.43%/0.484 -0.497/62.95%/0.306
LoRA (rank 4) 4-bit 0.109% -1.774/44.18%/0.474 -0.486/62.78%/0.291
LoRA (rank 8) 4-bit 0.217% -1.767/44.20%/0.484 -0.480/63.19%/0.296
LoRA (rank 16) 4-bit 0.434% -1.$7$0.475/62.70%/0.288
LoRA (rank 32) 4-bit 0.864% -1.771/44.39%/0.482 -0.475/63.24%/0.292
LoRA (rank 16) - 0.434% -1.774/44.15%/0.480 -0.475/62.84%/0.304
LoHa (rank 8) - 0.434% -1.762/44.23%/0.484 -0.483/62.91%/0.301
LoKr (rank 64) - 0.028% -1.762/44.17%/0.499 -0.470/63.13%/0.288
IA3 - 0.031% -1.770/43.99%/0.473 -0.487/63.24%/0.293
IA3 4-bit 0.031% -1.769/44.00%/0.477 -0.484/63.35%/0.296
(a) Log-likelihood
(b) Accuracy
(c) RMSE